OpenStudy (anonymous):

A family of 6 eats at an Italian restaurant. Some order a meatball sub for $9. Some order chicken parmesan for $11. The total bill was $58. How many ordered the meatball sub?

OpenStudy (anonymous):

4 meatball subs

OpenStudy (anonymous):

thank ya thank ya

OpenStudy (anonymous):

anytime finally I could help @kyle_bradley

OpenStudy (anonymous):

haha its ok

OpenStudy (abmon98):

9x+11(6-x)=58 9x+66-11x=58 -2x=-8 x=4 meatballs

OpenStudy (anonymous):

thanks @Abmon98

OpenStudy (abmon98):

Your most welcome :)